What is the oxidation number of Mn in the following coordination compound?

Li[Mn(NH3)2Br4]

[MnCl6]Cl
Answer #1

Oxidation number is an indicator of the degree of oxidation (loss of electrons) of an atom in a chemical compound.

Li[Mn(NH3)2Br4]

Li = +1

Mn = x

NH3 = 0

Br = -1

as the compound is neutral, we get

(+1) + x + (0) + (-1 * 4) = 0

+1 + x - 4 = 0

-3 + x = 0

x = +3

so the oxidation number of Mn in Li[Mn(NH3)2Br4] = +3

b) [MnCl6]Cl

Mn = x

Cl = -1

as compund is neutral we get

x + (-1)*6 + (-1) = 0

x - 6 - 1 = 0

x - 7 = 0

x = +7

so the oxidation number of Mn in [MnCl6]Cl is +7